A CAT 834H four-wheel tractor (rubber tires, high pressure), with an operating weight of 108,180 lb...

  1. A CAT 834H four-wheel tractor (rubber tires, high pressure), with an operating weight of 108,180 lb , is pulled up a road whose slope is + 3 .0% at a uniform speed.

  1. If the average tension in the towing cable is 6,060 lb, what is the rolling resistance of the road?

Hint:

  • The useable power for overcoming rolling resistance is equal to the tension in the towing table but some of it will be consumed on overcoming grade resistance
  • By having the usable tension power and the weight of the machine (in US tons) you may find the rolling resistance

  1. This would be what type of road surface?

Hint: Use table 6.1 to find the corresponding road surface for the rolling resistance you found.
Answer #1

Operating weight of tractor=(W)=108180 lb

Deduction= slope x weight of the tractor

=0.03x108180=3245.4 lb

Conversion of W to tonne:-

2000 lb=1 tonne

108180 lb=(108180/2000) =54.09 tonne

Tension required to overcome the rolling resistance=(P)=(6060-3245.4)lb=2814.6 lb

Rolling resistance(R)=P/W =(2814.6/54.09)=52.04 lb/ tonne

From table 6.1

The type of road surface corresponding to rolling resistance 52.09 lb/ tonne and Rubber tires with high pressure is= Earth,compacted and maintained surface.